HIGH-RESOLUTION SPECTROSCOPY OF THE 2^{2}Π_{u} ← X^{4}Σ_{g}^{−} FORBIDDEN TRANSITIONS OF C_{2}^{+}

\underline{C. G. TARSITANO}, C. F. NEESE, and T. OKA, Department of Chemistry, Department of Astronomy and Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

The electronic absorption spectrum of the (0,2), (1,3), and (6,9) bands of the B^{4}Σ_{u}^{−} − X^{4}Σ_{g}^{−} system of C_{2}^{+} was obtained using the velocity modulation technique in conjunction with heterodyne detection. The rotationally resolved spectrum shows perturbations, which are attributed to the 2^{2}Π_{u} state. The mixing between the B^{4}Σ_{u}^{−} state and the 2^{2}Π_{u} state for nearly degenerate levels generated enough intensity borrowing to observe twenty 2^{2}Π_{u} ← X^{4}Σ_{g}^{−} forbidden transitions. The bands, with their corresponding forbidden transitions, were fit to a model Hamiltonian. Line position measurements, line strength factors, and expectation values for the orbital angular momentum, ⟨Λ′⟩, for the forbidden transitions are reported. Molecular parameters from the global fit of each band, including their corresponding forbidden transitions, are reported.